The recipe for mint chocolate ice cream requires 2 and half cups of cream for 10 people.You need ice cream for 16 people.How much cream will you need?


1
Expert's answer
2022-01-10T16:29:22-0500
"\begin{matrix}
 2\dfrac{1}{2}\ cups & for & 10\ people \\
 x\ cups & for & 16\ people 
\end{matrix}"

"x=\dfrac{2\dfrac{1}{2}(16)}{10}=\dfrac{\dfrac{5}{2}(16)}{10}=4"



We will need 4 cups of cream for 16 people.